The base of an isosceles triangle is 12 in. less than the length of the equal sides. If the perimeter of the triangle is 39 in., find the length of each of the sides.
Answer by checkley75(3666)

B=L-12
L-12+L=L=39
3L=39+12
3L=51
L=51/3
L=17 IN. IS THE LENGTH OF THE SIDES.
BASE=17-5
BASE=12
PROOF
5+17+17=39
39=39